# Quantum-Enhanced Neural Exchange-Correlation Functionals

### v2

@@ VI Acknowledgements
The authors thank the PASQAL team for fruitful discussions, especially Atiyo Ghosh, Andrea A. Gentile and Louis-Paul Henry for their valuable insights.
@@ Appendix A Incorporating fractional occupation in RKS-DFT
Chai’s algorithm for fractional occupation is based on the argument that the distribution of occupation numbers is independent of the specific many-body system and adheres to a universal form described by the Fermi-Dirac distribution. Chai (2012) Hence, the density can be expressed as $n({\bf r})=\sum_{i=1}^{\infty}g_{i}|\psi_{i}({\bf r})|^{2},$ where the occupation number $g_{i}$ is the Fermi-Dirac function $g_{i}=\{1+\text{exp}[(\epsilon_{i}-\mu)/\gamma]\}^{-1},$ which obeys the following two conditions, $\sum_{i=1}^{\infty}g_{i}=N,$ and $0\leq g_{i}\leq 1,$ where $\epsilon_{i}$ is the orbital energy of the $i^{\text{th}}$ orbital $\psi_{i}({\bf r})$, and $\mu$ is the chemical potential chosen to maintain a constant number of electrons $N$. To account for this modification, the expression for the energy is then given by
[…]
@@
with $\gamma$-dependent XC energy given by $E^{\gamma}_{\text{XC}}[n]=E_{\text{XC}}[n]+E_{\gamma}[n]$, where $E_{\gamma}[n]$ represents the difference between the non-interacting kinetic free energy at zero temperature and that at temperature $\gamma$. The entropy contribution is given by $-\frac{\gamma}{k_{B}}S_{s}^{\gamma}\left[\left\{g_{i}\right\}\right]=\gamma\sum_{i=1}^{\infty}\left[g_{i}\ln\left(g_{i}\right)+\left(1-g_{i}\right)\ln\left(1-g_{i}\right)\right]$. Given a suitable electronic temperature $\gamma$, the chemical potential $\mu$ is fixed by $\sum_{i=1}^{\infty}\left\{1+\exp\left[\left(\epsilon_{i}-\mu\right)/\gamma\right]\right\}^{-1}=N$ to conserve the number of particles. In essence, only the entropy contribution needs to be added to an existing RKS-DFT implementation and $E^{\gamma}_{\text{XC}}[n]$ is learned by the neural XC. This approach does not impact the overall computational scaling. The trade-off for improved density representability is the inclusion of a single free parameter, the temperature $\gamma$, which may require initial tuning for consistency across systems, a subject that warrants further investigation.
@@ Appendix B Integration grids and neural XC functionals
Figure 9: Electron density of a hydrogen molecule ($R=3.84$ Bohr) at first six KS iterations using Global QNN within the 1D KS-DFT framework. At every KS iteration, the electron density is improved, eventually reaching the reference one. The reference (Ref) result is generated using DMRG by Li et al. Li et al. (2021).
A neural XC model needs to implicitly learn a function $f_{\theta}$ that is adapted to the $\{w_{i}\}^{N_{\mathrm{grid}}}_{i=1}$ grid points, distributed according to a fixed grid type. There are multiple choices of grid types and densities (number of grid points $N_{\mathrm{grid}}$), with a finer grid mesh yielding more accurate results at a higher computational cost. Ideally, a neural XC model needs to be trained/constrained to be invariant to finer mesh grids of the same type. See Fig. 9 where a trained Global QNN model is used to compute the electron density throughout KS iterations on a fixed grid of 513 points. This adds to the challenge of making generalizable neural XC functionals where grid (hence the number of model inputs) needs to be adapted and neural XC functional made consistent with the changes. For instance, a potential solution is to design a grid invariant embedding or add constraints in the loss to make the NN $f_{\theta}$ invariant to grid density changes. Another approach would be to use also the coordinates $\{\mathbf{r}_{i}\}_{i=1}^{N_{\mathrm{grid}}}$ as inputs of the model $f_{\theta}$ to explicitly include the information about the distribution of grid points in the Cartesian space. However, this approach would quadruple the number of inputs, so it is not currently considered for simplicity. In our work, we minimize grid variations by using a single grid type (e.g., of Lebedev typeLebedev (1975)) and the same number of grid points for all molecules whenever possible (our models have a fixed number of $N_{\mathrm{grid}}$ inputs) in order to facilitate the training.
@@ C.1 Quantum circuit ansatz
As an ansatz $U_{a}$ layer, we employ the popular Hardware-Efficient Ansatz (HEA) introduced in the context of VQE for chemistry. Peruzzo et al. (2014) The heuristic ansatz structure is not specifically tailored to the problem; thus, it represents the most general quantum model, closely resembling a classical MLP. It is composed of selected single qubit rotations (i.e., $R_{z}(\theta)=e^{-\frac{i}{2}\theta\hat{Z}}$) that are placed on all available qubits with distinct angles $\theta_{a}$,
@@ Appendix D Differentiation of quantum models
Gradient evaluation is an essential component for training NNs in general. Since (quantum) NNs can be thought of as a function $f_{\theta}$, there are various approaches for differentiation. The standard approach in classical ML is the Automatic Differentiation (AD) that exploits the chain rule and stores the intermediate values of subfunctions (layers) of ML models to evaluate the gradients of $f_{\theta}$. This approach has an advantageous scaling that enables training large classical models and, hence, is always used for the differentiation of classical layers in this work. In the context of quantum computing, there has been a need to develop an efficient approach, beyond finite differences, for gradient evaluation. Schuld et al. showed that due to the trigonometric relations within the unitaries associated with single qubit rotations, $U_{\mathrm{sqr}}=e^{-\frac{i}{2}\theta\hat{O}}$ (that include single-qubit Pauli generators $\hat{O}$), it is possible to derive a simple and exact derivative formula which highly resembles finite differences, the so-called Parameter-Shift Rule (PSR): Schuld et al. (2019)
@@
The generalized PSR (GPSR) rule Kyriienko and Elfving (2021) is needed for general multi-qubit generators.

### v3

@@ VI Acknowledgements
The authors thank the Pasqal team for fruitful discussions, especially Atiyo Ghosh, Andrea A. Gentile, John West and Louis-Paul Henry for their valuable insights.
@@ Appendix A Incorporating fractional occupation in RKS-DFT
Chai’s algorithm for fractional occupation is based on the argument that the distribution of occupation numbers is independent of the specific many-body system and adheres to a universal form described by the Fermi-Dirac distribution. frac-occ Hence, the density can be expressed as $n({\bf r})=\sum_{i=1}^{\infty}g_{i}|\psi_{i}({\bf r})|^{2},$ where the occupation number $g_{i}$ is the Fermi-Dirac function $g_{i}=\{1+\text{exp}[(\epsilon_{i}-\mu)/\gamma]\}^{-1},$ which obeys the following two conditions, $\sum_{i=1}^{\infty}g_{i}=N,$ and $0\leq g_{i}\leq 1,$ where $\epsilon_{i}$ is the orbital energy of the $i^{\text{th}}$ orbital $\psi_{i}({\bf r})$, and $\mu$ is the chemical potential chosen to maintain a constant number of electrons $N$. To account for this modification, the expression for the energy is then given by
$E[n]=$ $T_{s}^{\gamma}\left[\left\{g_{i},\psi_{i}\right\}\right]-\frac{\gamma}{k_{B}}S\left[\left\{g_{i}\right\}\right]+\int n(\mathbf{r})v(\mathbf{r})d\mathbf{r}$ (29)
@@
with $\gamma$-dependent XC energy given by $E^{\gamma}_{\text{XC}}[n]=E_{\text{XC}}[n]+E_{\gamma}[n]$, where $E_{\gamma}[n]$ represents the difference between the non-interacting kinetic free energy at zero temperature and that at temperature $\gamma$. The entropy contribution is given by $\frac{\gamma}{k_{B}}S_{s}^{\gamma}\left[\left\{g_{i}\right\}\right]=\gamma\sum_{i=1}^{\infty}\left[g_{i}\ln\left(g_{i}\right)+\left(1-g_{i}\right)\ln\left(1-g_{i}\right)\right]$. Given a suitable electronic temperature $\gamma$, the chemical potential $\mu$ is fixed by $\sum_{i=1}^{\infty}\left\{1+\exp\left[\left(\epsilon_{i}-\mu\right)/\gamma\right]\right\}^{-1}=N$ to conserve the number of particles. In essence, only the entropy contribution needs to be added to an existing RKS-DFT implementation and $E^{\gamma}_{\text{XC}}[n]$ is learned by the neural XC. This approach does not impact the overall computational scaling. The trade-off for improved density representability is the inclusion of a single free parameter, the temperature $\gamma$, which may require initial tuning for consistency across systems, a subject that warrants further investigation.
@@ Appendix B Integration grids and neural XC functionals
Figure 11: Electron density of a hydrogen molecule ($R=3.84$ Bohr) at first six KS iterations using Global QNN within the 1D KS-DFT framework. At every KS iteration, the electron density is improved, eventually reaching the reference one. The reference (Ref) result is generated using DMRG by Li et al. ksr.
A neural XC model needs to implicitly learn a function $f_{\theta}$ that is adapted to the $\{w_{i}\}^{N_{\mathrm{grid}}}_{i=1}$ grid points, distributed according to a fixed grid type. There are multiple choices of grid types and densities (number of grid points $N_{\mathrm{grid}}$), with a finer grid mesh yielding more accurate results at a higher computational cost. Ideally, a neural XC model needs to be trained/constrained to be invariant to finer mesh grids of the same type. See Fig. 11 where a trained Global QNN model is used to compute the electron density throughout KS iterations on a fixed grid of 513 points. This adds to the challenge of making generalizable neural XC functionals where grid (hence the number of model inputs) needs to be adapted and neural XC functional made consistent with the changes. For instance, a potential solution is to design a grid invariant embedding or add constraints in the loss to make the NN $f_{\theta}$ invariant to grid density changes. Another approach would be to use also the coordinates $\{\mathbf{r}_{i}\}_{i=1}^{N_{\mathrm{grid}}}$ as inputs of the model $f_{\theta}$ to explicitly include the information about the distribution of grid points in the Cartesian space. However, this approach would quadruple the number of inputs, so it is not currently considered for simplicity. In our work, we minimize grid variations by using a single grid type (e.g., of Lebedev typeLebedev-USSR-1975) and the same number of grid points for all molecules whenever possible (our models have a fixed number of $N_{\mathrm{grid}}$ inputs) in order to facilitate the training.
@@ C.1 Quantum circuit ansatz
As an ansatz $U_{a}$ layer, we employ the popular Hardware-Efficient Ansatz (HEA) introduced in the context of VQE for chemistry. peruzzo2014variational The heuristic ansatz structure is not specifically tailored to the problem; thus, it represents the most general quantum model, closely resembling a classical MLP. It is composed of selected single qubit rotations (i.e., $R_{z}(\theta)=e^{-\frac{i}{2}\theta\hat{Z}}$) that are placed on all available qubits with distinct angles $\theta_{a}$,
@@ Appendix D Differentiation of quantum models
Gradient evaluation is an essential component for training NNs in general. Since (quantum) NNs can be thought of as a function $f_{\theta}$, there are various approaches for differentiation. The standard approach in classical ML is the Automatic Differentiation (AD) that exploits the chain rule and stores the intermediate values of subfunctions (layers) of ML models to evaluate the gradients of $f_{\theta}$. This approach has an advantageous scaling that enables training large classical models and, hence, is always used for the differentiation of classical layers in this work. In the context of quantum computing, there has been a need to develop an efficient approach, beyond finite differences, for gradient evaluation. Schuld et al. showed that due to the trigonometric relations within the unitaries associated with single qubit rotations, $U_{\mathrm{sqr}}=e^{-\frac{i}{2}\theta\hat{O}}$ (that include single-qubit Pauli generators $\hat{O}$), it is possible to derive a simple and exact derivative formula which highly resembles finite differences, the so-called Parameter-Shift Rule (PSR): schuld2019evaluating
@@
The generalized PSR (GPSR) rule kyriienko2021generalized is needed for general multi-qubit generators.
Appendix E Number of measurements
To determine the measurement budget to reach an accuracy $\varepsilon$ with QNNs presented in this work, we follow the standard frequentist approach wecker2015progress; romero2018strategies. The output of QNNs is obtained by measuring the total magnetization cost operator, which can be expressed as the sum of the expectation values of Pauli-$Z$ operators over all $N_{{q}}$ qubits, $M=\sum_{i=1}^{N_{{q}}}\langle\hat{Z}_{i}\rangle.$ Assume each term $\hat{Z}_{i}$ is measured $m_{i}$ times. The statistical error on each term, $\varepsilon_{i}$, is given by $\varepsilon_{i}^{2}=\frac{\text{Var}[\langle\hat{Z}_{i}\rangle]}{m_{i}}.$ Since $\langle\hat{Z}_{i}\rangle\in[-1,1]$, its variance is upper bounded by 1, $\text{Var}[\langle\hat{Z}_{i}\rangle]\leq 1.$ To achieve a total error $\varepsilon$, we can distribute the per-term errors such that $\varepsilon_{i}^{2}=\frac{1}{N_{q}}\varepsilon^{2}.$ Then, the total number of measurements $N_{\mathrm{shots}}$ across all qubits becomes
$N_{\mathrm{shots}}=\sum_{i=1}^{N_{{q}}}m_{i}=\sum_{i=1}^{N_{{q}}}\frac{\text{Var}[\langle\hat{Z}_{i}\rangle]}{\varepsilon_{i}^{2}}\leq\sum_{i=1}^{N_{{q}}}\frac{1}{\varepsilon^{2}/N_{q}}=\frac{N_{{q}}^{2}}{\varepsilon^{2}}.$ (34)
Since all the operators are diagonal, they can be measured simultaneously, resulting in a total shot cost of $N_{\mathrm{shots}}=\mathcal{O}(\varepsilon^{-2})$.
Appendix F Noise channels
In Sec. E, we discussed the error arising from finite sampling. In modeling realistic quantum devices, additional noise processes are commonly represented by completely positive trace-preserving (CPTP) maps acting on a qubit’s density matrix $\rho=|\Psi\rangle\langle\Psi|$. Any CPTP map can be expressed in terms of a set of Kraus operators $\{\hat{K}_{i}\}$ satisfying $\sum_{i}\hat{K}_{i}^{\dagger}\hat{K}_{i}=\mathbb{I}$. Two common single-qubit noise channels are dephasing and amplitude damping, which we focused on in our experiment in Tab. 8 implemented in QEX qex2025pasqal.
Dephasing (phase-damping) channel. The dephasing channel describes the decay of quantum coherences without altering computational-basis populations. A Kraus representation is
$\hat{K}_{0}=\sqrt{1-p}\,\mathbb{I},\quad\hat{K}_{1}=\sqrt{p}\,\hat{Z}$, (35)
where $0\leq p\leq 1$ is the dephasing probability and $\hat{Z}$ is the Pauli-$Z$ operator. The channel acts as
$\mathcal{E}_{\mathrm{deph}}(\rho)=\hat{K}_{0}\rho\hat{K}_{0}^{\dagger}+\hat{K}_{1}\rho\hat{K}_{1}^{\dagger}.$ (36)
Amplitude-damping channel. The amplitude-damping channel models energy relaxation from $\lvert 1\rangle$ to $\lvert 0\rangle$, as in spontaneous emission. A standard Kraus representation is
$\hat{K}_{2}=\begin{pmatrix}1&0\\ 0&\sqrt{1-p}\end{pmatrix},\quad\hat{K}_{3}=\begin{pmatrix}0&\sqrt{p}\\ 0&0\end{pmatrix}$, (37)
where $p$ is the probability of decay during the noise process. The map is defined as
$\mathcal{E}_{\mathrm{amp}}(\rho)=\hat{K}_{2}\rho\hat{K}_{2}^{\dagger}+\hat{K}_{3}\rho\hat{K}_{3}^{\dagger}.$ (38)
In Sec. IV.7, we analyze the performance of a typical QNN in our framework by applying these noise channels and tuning their strength via the parameter $p$.
Appendix G Proofs of error bounds
Proof.
We prove Theorem 1 under assumptions 13, as stated in Sec. IV.7. We aim to show that, under an error $\varepsilon$, the exact self-consistent density $n^{\ast}$ obtained from the exact SCF map $\mathcal{T}$ (with the exact XC functional) differs from the approximate self-consistent density $\tilde{n}$ obtained from the approximate SCF map $\widetilde{\mathcal{T}}$ by at most $\mathcal{O}(\varepsilon)$.
Step 1: We define iteration errors. Let $e_{k}=\|\,n^{(k)}-n^{*}\|$, with $||\cdot||\equiv||\cdot||_{2}$. Then by definition of $\mathcal{T}$,
$n^{(k+1)}-n^{*}\;=\;\widetilde{\mathcal{T}}[n^{(k)}]\;-\;\mathcal{T}[n^{*}].$
We split and add $\mathcal{T}[n^{(k)}]$:
$n^{(k+1)}-n^{*}=\bigl(\widetilde{\mathcal{T}}-\mathcal{T}\bigr)\!\bigl[n^{(k)}\bigr]+\Bigl(\mathcal{T}[n^{(k)}]-\mathcal{T}[n^{*}]\Bigr).$
Hence,
$e_{k+1}$ $=\bigl\|\,n^{(k+1)}-n^{*}\bigr\|\leq\bigl\|\bigl(\widetilde{\mathcal{T}}-\mathcal{T}\bigr)[n^{(k)}]\bigr\|$
$\quad+\bigl\|\mathcal{T}[n^{(k)}]-\mathcal{T}[n^{*}]\bigr\|$
$\leq\underbrace{\|\widetilde{\mathcal{T}}[n^{(k)}]-\mathcal{T}[n^{(k)}]\|}_{\text{Term (a)}}+\underbrace{\|\mathcal{T}[n^{(k)}]-\mathcal{T}[n^{*}]\|}_{\text{Term (b)}}.$ (39)
Step 2: We bound Term (b) by contractivity. By Assumption 1, for $n_{1},n_{2}\in U$,
$\|\mathcal{T}[n_{1}]-\mathcal{T}[n_{2}]\|\;\leq\;\kappa\,\|n_{1}-n_{2}\|.$
Since $n^{(k)}$ and $n^{*}$ lie in $U$, we get
$\|\mathcal{T}[n^{(k)}]-\mathcal{T}[n^{*}]\|\;\leq\;\kappa\,\|\,n^{(k)}-n^{*}\|\;=\;\kappa\,e_{k}.$
Step 3: We bound Term (a) by uniform XC error. From Sec. IV.7, we recall that $\widetilde{\mathcal{T}}$ and $\mathcal{T}$ differ only in the XC potential part, and by Assumption 2, that potential difference is at most $K\,\varepsilon$. To convince yourself, see the KS equations in Sec. II, where Hamiltonian operator $\hat{H}[n]=-\nabla^{2}/2+v_{\mathrm{KS}}[n]$ ultimately defines the SCF map $\mathcal{T}$ through XC potential $v_{\mathrm{XC}}$. In the local region $U$, changing the potential by at most $K\,\varepsilon$ changes the resulting density by at most some constant factor $\alpha\,\varepsilon$. Hence we have
$\|\widetilde{\mathcal{T}}[n]-\mathcal{T}[n]\|\;\leq\;\alpha\,\varepsilon\quad\text{for all}\;n\in U.$ (40)
Thus, for $n^{(k)}\in U$,
$\|\widetilde{\mathcal{T}}[n^{(k)}]-\mathcal{T}[n^{(k)}]\|\;\leq\;\alpha\,\varepsilon.$
Step 4: We combine the previous statements into Eq. (39):
$e_{k+1}$ $\;=\;\|n^{(k+1)}-n^{*}\|$
$\;\leq\;\|\widetilde{\mathcal{T}}[n^{(k)}]-\mathcal{T}[n^{(k)}]\|+\|\mathcal{T}[n^{(k)}]-\mathcal{T}[n^{*}]\|$
$\;\leq\;\alpha\,\varepsilon\;+\;\kappa\,e_{k}.$ (41)
Step 5: The density $n^{(k)}$ remains in $U$ by Assumption 3, we solve the recurrence by rewriting
$e_{k+1}\;\leq\;\kappa\,e_{k}\;+\;\alpha\,\varepsilon,$
as
$e_{k}\;\leq\;\kappa^{k}\,e_{0}+\alpha\,\varepsilon\sum_{j=0}^{k-1}\kappa^{j}\;=\;\kappa^{k}\,e_{0}+\alpha\,\varepsilon\,\frac{\,1-\kappa^{k}\,}{\,1-\kappa\,}.$
Set $n^{(0)}$ as the initial guess, so $e_{0}=\|n^{(0)}-n^{*}\|$. This proves the finite-$k$ bound (27). Taking $k\to\infty$, we get
$e_{\infty}=\|\,\widetilde{n}-n^{*}\|\;\leq\;\frac{\alpha\,\varepsilon}{\,1-\kappa\,}.$ (42)
Hence the approximate self-consistent density $\widetilde{n}$ is within $\mathcal{O}(\varepsilon)$ of the true electron density $n^{*}$.
Step 6: Energy difference is also $O(\varepsilon)$. Finally, define $\widetilde{E}\equiv\widetilde{E}[\widetilde{n}$] and $E^{*}\equiv E[n^{*}]$. Write
$\widetilde{E}-E^{*}\;$ $=\;\bigl(\widetilde{E}[\widetilde{n}]-E[\widetilde{n}]\bigr)\;+\;\bigl(E[\widetilde{n}]-E[n^{*}]\bigr).$ (43)
Since $\widetilde{E}_{\mathrm{XC}}$ differs from $E_{\mathrm{XC}}$ by at most $\varepsilon$ in value, $\widetilde{E}$ differs from $E$ by at most $\varepsilon$ at the same density $\widetilde{n}$. Meanwhile, we assume that $|E[\widetilde{n}]-E[n^{*}]|\leq M\,\|\widetilde{n}-n^{*}\|$ for some Lipschitz constant $M>0$ near $n^{*}$ since according to Eq. (42), $\tilde{n}$ and $n^{\ast}$ are order $\varepsilon$ close in the region $U$ by Assumption 3. Hence,
$|\widetilde{E}-E^{*}|\leq\varepsilon+M\,\|\widetilde{n}-n^{*}\|\leq\varepsilon+M\,\frac{\alpha}{1-\kappa}\,\varepsilon=C^{\prime}\,\varepsilon,$
where $C^{\prime}=1+\frac{M\,\alpha}{1-\kappa}$. This completes the proof. ∎
Thus, the final approximate energy is $\mathcal{O}(\varepsilon)$-close to the true energy, demonstrating that, under our assumptions, catastrophic error accumulation in the SCF loop is avoided.
Appendix H Lipschitz continuity of QNNs
In this section, we establish the Lipschitz continuity of QNNs employed in this work, following the derivations and notation from ref. schuld2019evaluating for clarity. Recall the definition of quantum models, Eq. (13), here denoted as
$f_{\theta}(n)=\langle 0|U^{\dagger}(n,\theta)\,\hat{C}\,U(n,\theta)|0\rangle,$
which forms the basis of all our $N_{q}$-qubit models. We propose the following Lemma to support our arguments from Sec. IV.7.
Lemma 1 (QNN Lipschitz continuity).
Let QNN be represented by $f_{\theta}(n)=\langle 0\mid U^{\dagger}(n,\theta)\,\hat{C}\,U(n,\theta)\mid 0\rangle$ with
$U(n)$ $=W^{(L+1)}S(n)W^{(L)}\ldots W^{(2)}S(n)W^{(1)},$
$S(n)$ $=\;e^{-inH},\ \ \ \hat{C}=\sum_{i=1}^{N_{q}}\hat{Z}_{i},$
where $W^{k}\equiv W^{k}(\theta)$ are $x$-independent unitaries (representing ansatz), $S(n)$ is the feature map with Hermitian operator $H$ with eigenvalues $\{\lambda_{m}\}_{m=1}^{d}$, $d=2^{N_{q}}$, and $L=N_{l}$ is the number of layers. Then $f_{\theta}$ is globally Lipschitz in $x$, defined according to ref. schuld2019evaluating as
$f(n)=\sum_{\bm{k},\bm{j}\in[d]^{L}}e^{i\left(\Lambda_{\bm{k}}-\Lambda_{\bm{j}}\right)n}a_{\bm{k},\bm{j}},\qquad\Lambda_{j}=\lambda_{j_{1}}+\cdots+\lambda_{j_{L}},$
Consequently,
$\sup_{x\in\mathbb{R}}|f^{\prime}_{\theta}(n)|\leq N_{q}\sum_{\bm{k},\bm{j}\in[d]^{L}}|\Lambda_{\bm{k}}-\Lambda_{\bm{j}}|<\infty,$
where we use multi-index $\bm{j}=$ $\left\{j_{1},\ldots,j_{L}\right\}\in[d]^{L}$, $d]^{L}$ denotes any set of $L$ integers between $1$ and $d$, and $N_{q}$ is the number of qubits. Hence, $f_{\theta}$ is globally Lipschitz on $\mathbb{R}$ (for any density $n$) with a $\theta$-independent constant.
Proof.
We follow the ref. schuld2019evaluating and recall that:
$f(n)=\sum_{\bm{k},\bm{j}\in[d]^{L}}e^{i\left(\Lambda_{\bm{k}}-\Lambda_{\bm{j}}\right)n}a_{\bm{k},\bm{j}},\qquad$
where
$a_{\bm{k},\bm{j}}=\sum_{i,i^{\prime}}\left(W^{*}\right)_{1k_{1}}^{(1)}\left(W^{*}\right)_{j_{1}j_{2}}^{(2)}\ldots\left(W^{*}\right)_{j_{L}i}^{(L+1)}C_{i,i^{\prime}}$
$\times W_{i^{\prime}j_{L}}^{(L+1)}\ldots W_{j_{2}j_{1}}^{(2)}W_{j_{1}1}^{(1)}.$
Since $\|\hat{C}\|=N_{q}$, $W^{k}$ are unitary, and $\sup_{n}|f_{\theta}(n)|\leq N_{q}$, consequently the elements $|a_{\bm{k},\bm{j}}|\leq\|\hat{C}\|=N_{q}$. We can then write the derivative of our model,
$f_{\theta}^{\prime}(n)=\sum_{\bm{k},\bm{j}\in[d]^{L}}i\left(\Lambda_{\bm{k}}-\Lambda_{\bm{j}}\right)e^{i\left(\Lambda_{\bm{k}}-\Lambda_{\bm{j}}\right)n}a_{\bm{k},\bm{j}}.$
So $\sup_{n}|f^{\prime}_{\theta}(n)|\leq\sum_{\bm{k},\bm{j}\in[d]^{L}}|\Lambda_{\bm{k}}-\Lambda_{\bm{j}}|\,|a_{\bm{k},\bm{j}}|\leq N_{q}\sum_{\bm{k},\bm{j}\in[d]^{L}}|\Lambda_{\bm{k}}-\Lambda_{\bm{j}}|<\infty$ since the spectrum is finite. The mean‑value theorem implies the stated Lipschitz continuity of $f_{\theta}(n)$. ∎